HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

MID-INFRARED SATURATION SPECTROSCOPY FOR VIBRATIONAL FREQUENCY MEASURMENTS AT THE SUB-KHZ LEVEL

\underline{Y. LIU}, N. CAHUZAC, E. CANTIN, O. LOPEZ, D.B.A. TRAN, R. SANTAGATA, A. AMY-KLEIN, M. MANCEAU, B. DARQUIE, Laboratoire de Physique des Lasers, Université Sorbonne Paris Nord, CNRS, Villetaneuse, France; M. TONNES, B. POINTARD, M. ABGRALL, L. LORINI, Y. LE COQ, R. LE

TARGAT, D. XU, P.-E. POTTIE, LNE-SYRTE, Observatoire de Paris, PSL Research University, CNRS, Sorbonne Universités, UPMC, Paris, France; H. ALVAREZ-MARTINEZ, LNE-SYRTE, Observatoire de Paris, PSL Research University, CNRS, Sorbonne Universités, UPMC, Paris, France; Real Instituto y

Observatorio de la Armada (ROA), Cádiz, Spain

Molecular systems, owing to their numerous degrees of freedom, offer promising perspectives for improving tests of fundamental physics and precision measurements in general^{1}^{,}^{2}^{,}^{3}^{,}^{4}^{,}^{5}. In our group at Laboratoire de Physique des Laser, using an optical frequency comb, a 10 μm quantum cascade laser (QCL) is stabilized at the sub-Hz level to an ultra-stable near infrared reference signal operated at the French metrology institute. This signal is calibrated there to some of the best atomic clocks and transferred through a noise-compensated 26-km long fiber cable of the REFIMEVE infrastructure^{6}. This results in a record relative frequency uncertainty of 10^{−14} and a 0.1 Hz QCL linewidth^{7}.

We have used this ultra-stable and frequency controlled mid-infrared source to perform high resolution spectroscopy on various molecular species of fundamental, atmospheric and astrophysics interest such as methanol, ammonia or trioxane, with a few kilohertz uncertainties on line positions using sub-Doppler saturated absorption spectroscopy in a multipass-cell^{8}.

Here, I will discuss our latest efforts in order to improve these spectroscopic measurements, by realizing cavity enhanced measurements. I will show results on methanol achieving a few hundreds of Hz uncertainties on molecular frequencies, I will then present investigations on various sources of noise currently limiting our signal-to-noise ratio and perspectives for improvements. These results enhance our ability to meet the needs for our ongoing efforts towards studying the variation of fundamental constants^{3} or testing fundamental symmetries, such as the measurement of the tiny parity-violating energy difference between enantiomers of a chiral molecule^{1,2}, via precise molecular spectroscopy. Furthermore, it can contribute to the enrichment and refinement of the HITRAN database^{9}, which can be highly beneficial for the fields of astronomy^{10} and atmospheric physics^{11}, providing valuable data for accurate modeling and analysis.

1 doi:10.1070/QEL16880, A. Cournol et al., Quantum Electronics 49, 288 (2019)

2 doi:10.1021/acs.jpclett.2c02434, M.R. Fiechter et al., J. Phys. Chem. Lett. 13, 42, (2022)

3 doi:10.1140/epjqt/s40507-022-00130-5, G. Barontini et al., Quantum Technol. 9, 12, (2022)

4doi:10.1103/PhysRevLett.127.043201, J. Lukusa Mudiayi et al., Phys. Rev. Lett. 127, 043201, (2021)

5doi:10.1088/1681-7575/aaa790, J. Fischer et al., Metrologia. 55, R1 (2018)

6www.refimeve.fr, REseau FIbré MEtrologique à Vocation Européenne

7doi:10.1038/nphoton.2015.93, B. Argence et al., Nature Photon. 9, 456-460 (2015)

8doi:10.1364/OPTICA.6.000411, R. Santagata et al., Optica 6, 411 (2019)

p-number: p177  Submitted on Thu, 15 Jun 2023 13:49:54 +0200

9doi:10.1016/j.jqsrt.2021.107949, I. Gordon et al., J. Quant. Spectrosc. Radiat. Transfer 277, 107949 (2022)

10doi:10.1038/326049a0, W. Batrla et al., Nature 326, 49-51 (1987)

11doi:10.1029/98JD02747, D. Simpson et al., J. Geophys. Res. 104 (D7), 8113-8152 (1999)

p-number: p177  Submitted on Thu, 15 Jun 2023 13:49:54 +0200